OpenStudy (anonymous):

how many solutions do these equations have y=2x+5 y=-0.2x-4

OpenStudy (dumbsearch2):

Have you done the first step?

OpenStudy (dumbsearch2):

To start, find the y value at the place they intersect, therefore as y=2x + 5 and y=-0.2x -4 they intersect when the right hand of the equations are equal to each other.

terenzreignz (terenzreignz):

Or you could just note that they have different slopes. :D

OpenStudy (anonymous):

To find solutions to equations: Equate both solutions together: y=2x+5 y=-0.2x-4 i.e. 2x + 5 = -0.2x - 4 Make x the subject of the equation, and solve for x. Solve for y after by subbing the value of x into any of those y = equations.

OpenStudy (anonymous):

both of the equations are STRAIGHT LINES how many possible ways could 2 straight lines cross? Clearly only 1, unless both lines were exactly the same (which in this case they aren't)